OpenStudy (anonymous):

Calculate the average rate of change for the given graph from x = –2 to x = 0 and select the correct answer below.

OpenStudy (anonymous):

Courtesy of Texas Instruments 0 1 –1 –2

OpenStudy (anonymous):

@Nnesha @nincompoop @ganeshie8

OpenStudy (anonymous):

@Baby_Bear69

OpenStudy (baby_bear69):

idk , i'd be guessing C

OpenStudy (anonymous):

C is always my first guess

OpenStudy (anonymous):

@Jaynator495

OpenStudy (anonymous):

@paki

OpenStudy (baby_bear69):

@paki

OpenStudy (baby_bear69):

Haha we tagged him at the same time

OpenStudy (anonymous):

Do you know the formula for the average rate of change?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[M = \frac{ {y2}-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

so -1-(-1) 0-2

OpenStudy (anonymous):

2,-2

OpenStudy (anonymous):

or just -2

OpenStudy (anonymous):

So they're asking for the average rate of change from x = -2 to x = 0. So we have two ordered pairs. (-2, y1) and (0, y2). On the graph it shows that the y's are both -1. So yes you plug in the ordered pairs into the formula. \[M = \frac{ -1 - (-1) }{ 0 - (-2) }\]

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

What part are you confused on? I'll try to clear it up. :)

OpenStudy (anonymous):

when i solved the formula i got 2 and that isent one of the answer choices

OpenStudy (anonymous):

You shouldn't get an ordered pair when you solve it. Also, when you put -1 - (-1) 0-2. You forgot that the 2 is also negative so it'd actually be 0 - (-2). The subtraction signs become addition/positive of course. So your equation becomes \[M = \frac{ -1+1 }{ 0 + 2 }\]